# 4He monolayer on Graphene: a Quantum Monte Carlo study

### v2

@@ 1. Introduction
An interesting theoretical question is whether a phase simultaneously displaying diagonal (i.e., crystalline) and off-diagonal (i.e., superfluid) long-range order may occur at sufficiently low temperature, in some well-defined range of 4He coverage. Some experimental claims have been made of concurrent superfluidity and crystalline order in the second layer of 4He adsorbed on graphite, specifically at, or in the vicinity, of a solid phase registered with the underlying first (solid) layer [5, 6, 7, 8]. Though the denomination supersolid is, strictly speaking, not applicable to a system of this type [9], the observation of such a phase would nonetheless be of great significance, as it would help elucidate the interplay between these two seemingly antithetic types of order. It seems fair to state, however, that at this time all the evidence is indirect and/or inconclusive; indeed, the interpretation of the findings has been questioned and is not supported by reliable microscopic theoretical calculations [10, 11, 12, 13, 14].
Graphene, which consists of a sheet of graphite, is also regarded as a potentially interesting adsorber. A single layer of carbon atoms is about 10% less attractive than graphite; the ensuing enhancement of the quantum excursions of helium atoms away from the basal plane acts to soften the repulsive interaction among them at short distances. All of this may ultimately increase the mobility of 4He, possibly stabilizing novel phases displaying a more quantal character (e.g., superfluid phases with an unusual degree of atomic localization). Early microscopic calculations for helium on graphite showed that a mere 10% reduction in the strength of the attractive part of the interaction potential results in no significant change in the overall physical behavior of an adsorbed helium film [10], a conclusion confirmed by subsequent calculations for 4He on graphene. In particular, there is general agreement on the existence of the same commensurate crystalline phase of a 4He monolayer (known as $C_{1/3}$) which forms on graphite. There is, however, some controversy concerning a possible finite superfluid response of such a crystalline phase, at sufficiently low temperature, as well as on the existence of superfluid phases at lower coverage, either liquid or crystalline (doped with vacancies).
@@ 2. Model and methodology
We consider an assembly of $N$ 4He atoms, regarded as point-like identical particles of mass $m$ and spin $S=0$, moving in three physical dimensions. The system is enclosed in a cell of volume $\Omega=L_{x}\times L_{y}\times L_{z}$, with $L_{x}=34.08$, $L_{y}=29.514$, and $L_{z}=40$. Periodic boundary conditions are used in the three directions (but the boundary condition along the $z$ direction is immaterial, for reasons that will soon be clear). The graphene sheet is modeled as an ideal, 2D (honeycomb) lattice; our simulation cell accommodates $N_{c}=384$ carbon atoms at fixed positions on the $z=0$ plane, with a carbon-carbon bond length $a=1.42$. Regarding the carbon atoms as pinned at classical lattice sites is an assumption that has been made in all other comparable studies, and one that seems justified by their relatively large mass, compared to that of the helium atoms. The 4He coverage (i.e., 2D density) is $\theta=N/A$, where $A=L_{x}\times L_{y}$.
@@ 3.1. Finite-temperature results
Figure 1: Energy per 4He atom (in K) computed by simulation as a function of the temperature at a coverage $\theta=0.0636$-2. Statistical errors are smaller than symbol sizes. The solid line represents a fit to the data obtained with the expression $e(T)=e_{0}+\alpha T^{3}$. The estimate shown with a diamond at $T=0$ is that obtained with PIGS (see section 3.2) for a projection time $\Lambda=$ 1 K-1. When not shown, statistical errors are smaller than symbol sizes.
The data can be fitted with the expression $e(T)=e_{0}+\alpha T^{3}$, yielding $e_{0}=-129.550(25)$ K. The first observation is that energy estimates at finite temperature (for $T\leq 1$ K), for the particular coverage considered here are significantly lower (by about 0.25 K) than the (supposedly “exact”) ground state value quoted in Ref. [18], based on a calculation making use of the same potentials utilized in this work [29].
@@ 3.1.1. Low coverage results
It is contended in Ref. [15] that, at coverage less than $\theta_{0}$, permutations of identical helium atoms begin to occur, leading at sufficiently low temperature to a finite superfluid response, which for sufficiently low coverage is interpreted as the stabilization of a superfluid liquid-like state. In our simulations we also observe such permutations, in fact even at temperatures as high as 1 K (as shown in Fig. 3). They typically involve atoms that lack some nearest neighbors, and in a small simulation cell cycles involving several particles occasionally appear, connecting two opposite sides of the cell and resulting in a significant superfluid signal, which, for relatively small system sizes (e.g., $\sim$ 30 atoms) may appear as a bulk superfluid response. This is a finite-size effect, however; in the thermodynamic limit coexistence of a crystal with a low-density vapor is expected. Altogether, therefore, although the anisotropic He-C potential has a stronger pinning effect on the helium atoms, nonetheless the basic physics of the system below commensurate coverage is the same, regardless of whether an isotropic or an anisotropic potential is used. The physical picture that emerges from our simulations below commensurate coverage is the formation of solid 4He clusters, analogously to what is observed on graphite [36].
@@ 3.2. Ground state results
The study of the ground state of the system described by Eq. 1 was carried out using the following trial wave function:
@@
Although the expression (6) aims at providing a good fit to the data only in the $t\to\Lambda$ limit, it turns out to be possible to obtain an acceptable fit, essentially over the whole imaginary time range, as shown by the left side of Fig. 6. The probability distribution for the value of the parameter $\rho_{S}$ of the fit, shown on the left side of Fig. 6 places our estimate of the superfluid fraction in the ground state at less than $\sim 0.005$ with 99.99% confidence, which is already well over an order of magnitude lower than the estimate of Ref. [17]. We repeated this analysis for different values of $\Lambda$ but find a lower bound for $\rho_{S}$ greater than zero. Thus, we conclude that our results are consistent with a null value of $\rho_{S}$ in the ground state, in accord with all finite temperature studies.

### v3

@@ 1. Introduction
An interesting theoretical question is whether a phase simultaneously displaying diagonal (i.e., crystalline) and off-diagonal (i.e., superfluid) long-range order may occur at sufficiently low temperature, in some well-defined range of 4He coverage. Some experimental claims have been made of concurrent superfluidity and crystalline order in the second layer of 4He adsorbed on graphite, specifically at, or in the vicinity, of a solid phase registered with the underlying first (solid) layer [5, 6, 7, 8]. Though the denomination supersolid is, strictly speaking, not applicable to a system of this type [9, 10], the observation of such a phase would nonetheless be of great significance, as it would help elucidate the interplay between these two seemingly antithetic types of order. It seems fair to state, however, that at this time all the evidence is indirect and/or inconclusive; indeed, the interpretation of the findings has been questioned and is not supported by reliable microscopic theoretical calculations [11, 12, 13, 14, 15].
Graphene, which consists of a sheet of graphite, is also regarded as a potentially interesting adsorber. A single layer of carbon atoms is about 10% less attractive than graphite; the ensuing enhancement of the quantum excursions of helium atoms away from the basal plane acts to soften the repulsive interaction among them at short distances. All of this may ultimately increase the mobility of 4He, possibly stabilizing novel phases displaying a more quantal character (e.g., superfluid phases with an unusual degree of atomic localization). Early microscopic calculations for helium on graphite showed that a mere 10% reduction in the strength of the attractive part of the interaction potential results in no significant change in the overall physical behavior of an adsorbed helium film [11], a conclusion confirmed by subsequent calculations for 4He on graphene. In particular, there is general agreement on the existence of the same commensurate crystalline phase of a 4He monolayer (known as $C_{1/3}$) which forms on graphite. Indeed, the basic features of the phase diagram of an adsorbed 4He monolayer on graphene can be understood on the basis of a highly simplified description of the system, making use of a lattice Hamiltonian of hard core bosons [16].
There is, however, some controversy concerning a possible finite superfluid response of such a crystalline phase, at sufficiently low temperature, as well as on the existence of superfluid phases at lower coverage, either liquid or crystalline (doped with vacancies).
@@ 2. Model and methodology
We consider an assembly of $N$ 4He atoms, regarded as point-like identical particles of mass $m$ and spin $S=0$, moving in three physical dimensions. The system is enclosed in a cell of volume $\Omega=L_{x}\times L_{y}\times L_{z}$, with $L_{x}=34.08$, $L_{y}=29.514$, and $L_{z}=40$. Periodic boundary conditions are used in the three directions (but the boundary condition along the $z$ direction is immaterial, as helium atoms remain confined in the vicinity of the substrate). The cell sizes are chosen to accommodate an integer number of graphene unit cells in both directions, without introducing frustration. The graphene sheet is modeled as an ideal, 2D (honeycomb) lattice; our simulation cell accommodates $N_{c}=384$ carbon atoms at fixed positions on the $z=0$ plane, with a carbon-carbon bond length $a=1.42$. Regarding the carbon atoms as pinned at classical lattice sites is an assumption that has been made in all other comparable studies, and one that seems justified by their relatively large mass, compared to that of the helium atoms. The 4He coverage (i.e., 2D density) is $\theta=N/A$, where $A=L_{x}\times L_{y}$.
@@ 3.1. Finite-temperature results
Figure 1: Energy per 4He atom (in K) computed by simulation as a function of the temperature at a coverage $\theta=0.0636$-2. Statistical errors are smaller than symbol sizes. The solid line represents a fit to the data obtained with the expression $e(T)=e_{0}+\alpha T^{3}$. The estimate shown with a diamond at $T=0$ is that obtained with PIGS (see section 3.2) for a projection time $\Lambda=$ 1 K-1, while the black box refers to the estimate of Ref. [20]. When not shown, statistical errors are smaller than symbol sizes.
The data can be fitted with the expression $e(T)=e_{0}+\alpha T^{3}$, yielding $e_{0}=-129.550(25)$ K. The first observation is that energy estimates at finite temperature (for $T\leq 1$ K), for the particular coverage considered here are significantly lower (by about 0.25 K) than the ground state estimate quoted in Ref. [20], based on a calculation making use of the same potentials utilized in this work [31].
@@ 3.1.1. Low coverage results
It is contended in Ref. [17] that, at coverage less than $\theta_{0}$, permutations of identical helium atoms begin to occur, leading at sufficiently low temperature to a finite superfluid response, which for sufficiently low coverage is interpreted as the stabilization of a superfluid liquid-like state. In our simulations we also observe such permutations, in fact even at temperatures as high as 1 K (as shown in Fig. 3). They typically involve atoms that lack some nearest neighbors, and in a small simulation cell cycles involving several particles occasionally appear, connecting two opposite sides of the cell and resulting in a significant superfluid signal, which, for relatively small system sizes (e.g., $\sim$ 30 atoms) may appear as a bulk superfluid response. This is a finite-size effect, however; in the thermodynamic limit coexistence of a crystal with a low-density vapor is expected. In all of our simulations, the estimate of the superfluid fraction obtained as an average over sufficiently long runs is zero within statistical errors (i.e., the average value is typically much smaller than the statistical error). This is true both at commensurate and at lower coverage.
Altogether, therefore, although the anisotropic He-C potential has a stronger pinning effect on the helium atoms, nonetheless the basic physics of the system below commensurate coverage is the same, regardless of whether an isotropic or an anisotropic potential is used. The physical picture that emerges from our simulations below commensurate coverage is the formation of solid 4He clusters, analogously to what is observed on graphite [38].
@@ 3.2. Ground state results
In the case of ground state simulations, we restricted ourselves to the case of commensurate coverage $\theta_{0}$. The study of the ground state of the system described by Eq. 1 was carried out using the following trial wave function:
@@
Although the expression (6) aims at providing a good fit to the data only in the $t\to\Lambda$ limit, it turns out to be possible to obtain an acceptable fit, essentially over the whole imaginary time range, as shown by the left side of Fig. 6. The probability distribution for the value of the parameter $\rho_{S}$ of the fit, shown on the left side of Fig. 6 places our estimate of the superfluid fraction in the ground state at less than $\sim 0.005$ with 99.99% confidence, which is already well over an order of magnitude lower than the estimate of Ref. [19]. We repeated this analysis for different values of $\Lambda$ but we never obtained a lower bound for $\rho_{S}$ greater than zero, with an acceptable statistical confidence. Thus, we conclude that our results are consistent with a null value of $\rho_{S}$ in the ground state, in accord with all finite temperature studies. Because this conclusion is established on a finite system, a fortiori it must hold the thermodynamic limit.